# Optimal low-rank approximations for linear Gaussian inverse problems on Hilbert spaces Part II: posterior mean approximation

### v2

@@ 3 Equivalent Gaussian measures and Bayesian inference
We quantify posterior approximation errors using various divergences. Let $\nu_{2}$ be a target measure on $\mathcal{H}$ and $\nu_{1}$ an approximation of $\nu_{2}$ satisfying $\nu_{1}\sim\nu_{2}$. We use the $\rho$-Rényi divergence, the forward KL divergence, the Amari $\alpha$-divergence for $\alpha\in(0,1)$ and the Hellinger distance, defined respectively by,
@@
We refer to $D_{\textup{KL}}(\nu_{1}\|\nu_{2})$ as the ‘reverse KL divergence’. We do not distinguish between forward Rényi divergences $D_{\textup{Ren},\rho}(\nu_{2}\|\nu_{1})$ and reverse Rényi divergences $D_{\textup{Ren},\rho}(\nu_{1}\|\nu_{2})$, because of the ‘skew symmetry’ of the Rényi divergence: $D_{\textup{Ren},\rho}(\nu_{1}\|\nu_{2})=D_{\textup{Ren},1-\rho}(\nu_{2}\|\nu_{1})$, c.f. [36, Proposition 2].
@@
where (6) follows by [23, eqs. (134)–(135)]. It is then straightforward to show, c.f. [6, Remarks 3.10 and 3.11] that minimising the Amari-$\alpha$ divergence over $\nu_{1}$ is equivalent to minimising the $\alpha$-Rényi divergence over $\nu_{1}$. Furthermore, minimising the Hellinger distance over $\nu_{1}$ is equivalent to minimising the $\frac{1}{2}$-Rényi divergence over $\nu_{1}$. The divergence $D_{\textup{Ren},\frac{1}{2}}$ is also known as the Bhattacharyya distance, and is a metric.
@@
For a proof, see e.g. [3, Corollary 6.4.11] or [12, Theorem 2.25]. For injective covariances $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ such that items (i) and (iii) in Theorem 3.2 hold, we define
@@
If $\nu_{1}$ and $\nu_{2}$ are Gaussian measures, then the above divergences can be expressed explicitly in terms of the means and covariances of $\nu_{1}$ and $\nu_{2}$ using $R(\cdot\|\cdot)$ defined in (7). These formulations rely on a generalisation of the determinant to infinite-dimensional Hilbert spaces. For $A\in L_{1}(\mathcal{H})$, the so-called ‘Fredholm determinant’ $\det(I+A)$ can be defined, and if only $A\in L_{2}(\mathcal{H})$, then the notion of ‘Hilbert–Carleman determinant’ $\det_{2}(I+A)$ can be used. The Fredholm and Hilbert–Carleman determinants are defined on respectively trace-class and Hilbert–Schmidt perturbations of the identity. In finite dimensions, every operator is a trace-class and Hilbert–Schmidt perturbation of the identity, and hence these generalised determinants are defined everywhere in this case. In fact, the Fredholm determinant agrees with the standard determinant in this case. We refer to [29, Theorem 3.2, Theorem 6.2] or [30, Lemma 3.3, Theorem 9.2] for details.
The following result holds when $\mathcal{H}$ is a separable Hilbert space of finite or infinite dimension. The proof is a direct application of [23, Theorems 14 and 15].
@@
[6, Theorem 3.8] Let $m_{1},m_{2}\in\mathcal{H}$ and $\mathcal{C}_{1},\mathcal{C}_{2}\in L_{2}(\mathcal{H})_{\mathbb{R}}$ be positive. If $\mathcal{N}(m_{1},\mathcal{C}_{1})\sim\mathcal{N}(m_{2},\mathcal{C}_{2})$, then
@@
The prior and posterior distributions in (1) are equivalent, for every realisation $y$ in a set of probability 1, c.f. [34, Theorem 6.23]. The Hessian $H$ defined in (2) has rank $n$, hence the posterior precision is a finite-rank update of the prior by (3c). Using the operators $R(\mathcal{C}_{\textup{pr}}\|\mathcal{C}_{\textup{pos}})$ and $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ and Theorem 3.2, we can obtain the following relations between the prior-preconditioned Hessian $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}$ in (9a), the posterior-preconditioned Hessian in (9b), and the ‘pencil’ defined by the prior and the posterior covariance in (9c). Recall the notation $v\otimes w$ for $u,w\in\mathcal{H}$ from Section 1.4.
@@
[6, Proposition 3.7] There exists a nondecreasing sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ consisting of exactly $\operatorname{rank}\left({H}\right)$ nonzero elements and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for every $i\in\mathbb{N}$, and
@@
We note that the eigenvalues $(\frac{-\lambda_{i}}{1+\lambda_{i}})_{i}$ of (9a) relate to the eigenvalues $(\delta_{i}^{2})_{i}$ of [33, eq. (2.8)] via the transformation $\lambda_{i}=\eta(\delta_{i}^{2})$, $\delta_{i}^{2}=\eta(\lambda_{i})$ with $\eta(x)=\frac{-x}{1+x}$ for $x\in(-1,\infty)$.
@@
The significance of (10) is that the posterior variance along the span of $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ is smaller than the prior variance along the same subspace by a factor of $(1+\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{-1}$, for $i\in\mathbb{N}$. This was observed in the finite-dimensional case in [33, eq. (3.4)]. Thus, Proposition 3.4 implies that finite-dimensional data can only inform finitely many directions in parameter space, in the sense that posterior variance is reduced relative to prior variance only over a finite-dimensional subspace. The directions $(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})_{i\leq\operatorname{rank}\left({H}\right)}$ are orthogonal with respect to the $\mathcal{C}_{\textup{pr}}$-weighted inner product $\langle h_{1},h_{2}\rangle_{\mathcal{C}_{\textup{pr}}}\coloneqq\langle\mathcal{C}_{\textup{pr}}h_{1},h_{2}\rangle$, and not the unweighted inner product of $\mathcal{H}$.
@@
are subject to the largest relative variance reduction possible. Since $\mathcal{C}_{\textup{pr}}^{1/2}$ is injective, we thus conclude the following: among all $r$-dimensional subspaces of $\mathcal{H}$, it is the $r$-dimensional subspace $W_{r}$ that contains those $r$ linearly-independent directions in which the relative variance reduction is largest. This generalises the conclusion of [33, Section 3.1] to infinite dimensions.
@@ 4 Optimal approximation of the covariance
This section discusses low-rank posterior covariance approximation, using [6, Theorem 4.21]. This approximation serves as a basis for the joint mean and covariance approximation discussed in Section 6.
@@
Since $\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathscr{C}_{r}$ is positive and self-adjoint, it is an injective covariance operator. Furthermore, it is stated in [6, Corollary 4.9] that for every $\mathcal{C}\in\mathscr{C}_{r}$ it holds that $\mathcal{N}(m_{\textup{pos}}(y),\mathcal{C})$ is equivalent to the exact posterior. Since $\mathcal{C}_{\textup{pos}}$ does not depend on $y$, this equivalence holds for all $y$ simultaneously. This equivalence holds because of the range condition $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. Furthermore, the assumption $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ implies the rank restriction $\operatorname{rank}\left({K}\right)\leq r$. Thus, for $r$ small compared to $n$, $\mathcal{C}_{\textup{pr}}-KK^{*}$ can be interpreted as a low-rank update of $\mathcal{C}_{\textup{pr}}.$ Therefore, the class $\mathscr{C}_{r}$ provides an extension to infinite dimensions of the finite-dimensional updates considered in [33].
@@
The solution to Problem 4.1 now follows directly from [6, Lemma 4.2(iii)] and from [6, Theorem 4.21] applied with $f(x)\leftarrow f_{\textup{KL}}(\tfrac{-x}{1+x})$ , where
@@
Note that the results of [6, Theorem 4.21 and Corollary 4.23] are more general. They state that $\mathcal{C}^{\textup{opt}}_{r}$ is not only the optimal low-rank approximation of $\mathcal{C}_{\textup{pos}}$ for the reverse KL divergence, but simultaneously also for all divergences in a more general class of divergences, including the forward KL divergence, the Hellinger distance, the Rényi divergences and the Amari $\alpha$-divergences for $\alpha\in(0,1)$.
@@ 5 Optimal approximation of the mean
We shall assess the approximation quality of an approximate posterior mean by averaging the mean-dependent term for the Rényi divergence and the forward and reverse KL divergence over all possible realisations $y$ of $Y$. By averaging over $Y$, the optimal operator $A$ will be data-independent, i.e. will not depend on a specific realisation $y$ of $Y$. While averaging over $Y$ implies that the resulting posterior mean approximations are not optimal in general for a specific realisation $y$ of $Y$, this approach has the benefit that $A$ can be constructed before observing the data. This leads to an offline-online approach to posterior mean approximation: the preliminary ‘offline’ stage computes one operator, which can then can be applied in the subsequent ‘online’ stage to any realisation of the data. This is in analogy to the finite-dimensional case studied in [33, Section 4.1] and its generalisation to certain nonlinear forward models and to losses with respect to the average Amari $\alpha$-divergences as studied in [22, Section 5]. Furthermore, averaging over $Y$ enables us to exploit recent work on reduced-rank operator approximation [5].
@@
The Bayes risk $\mathcal{R}(A)\coloneqq\mathbb{E}\left[\left\lVert AY-X\right\rVert_{\mathcal{C}_{\textup{pos}}^{-1}}^{2}\right]$ for $A\in\mathscr{M}_{r}^{(i)}$, $i=1,2$, considered in [33, Section 4.1] is not well-defined, since the event $\{X\in\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{-1/2}}\}$ occurs with probability 0. However, one can show that $\mathcal{R}(A)=\mathbb{E}\left[\left\lVert AY-m_{\textup{pos}}(Y)\right\rVert_{\mathcal{C}_{\textup{pos}}^{-1}}^{2}\right]+\dim{\mathcal{H}}$ if $\dim{\mathcal{H}}<\infty$. Thus, not only does the approximation error (18) used in Problem 5.1 have a natural interpretation as the mean-dependent term of the Rényi, Amari, forward and reverse KL divergences, it also captures the relevant contribution to the Bayes risk which involves the approximation.
@@
~$\begin{split}\widetilde{\mathscr{M}}^{(1)}_{r}&\coloneqq\{(S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}-\widetilde{B})G^{*}\mathcal{C}_{\textup{obs}}^{-1}:\ \widetilde{B}\in\mathcal{B}_{00,r}(\mathcal{H})\},\\ \widetilde{\mathscr{M}}^{(2)}_{r}&\coloneqq\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H}).\end{split}$ (23)
@@
It is shown in items (iii) and (iv) of the following result that Problem 5.4 is a reformulation of Problem 5.1. Using Theorem 3.2, item (i) of the following result also provides an explicit description of the approximation classes $\mathscr{M}^{(i)}_{r}$ of (4) in terms of the ranges of the operators $A$ and $B$, while item (ii) relates these classes to the classes ~$\widetilde{\mathscr{M}}^{(i)}_{r}$ from (23).
@@
Let $r\leq n$ and $i=1,2$. Let $S_{\textup{pos}}$ be as defined in (21), let $\mathscr{M}^{(i)}_{r}$ be as in (4) and let ~$\widetilde{\mathscr{M}}^{{(i)}}_{r}$ be as in (23). Then,
@@
    ~$\widetilde{\mathscr{M}}^{(i)}_{r}=S_{\textup{pos}}^{-1}\mathscr{M}^{(i)}_{r}$,
@@
In order to solve Problem 5.4, we use a result on reduced-rank operator approximation in $L_{2}(\mathcal{H})$ norm, proven in [5]. It is a generalised version of the Eckhart–Young theorem. Recall that compact operators, in particular Hilbert–Schmidt operators and finite-rank operators, have an SVD, c.f. Lemma A.3. Also recall the definition of the Moore–Penrose inverse $C^{\dagger}$ of $C\in\mathcal{B}(\mathcal{H})$ from Section 1.4. If $C$ has closed range, then $C^{\dagger}$ is bounded, c.f. [13, Proposition 2.4]. The following is an application of [5, Theorem 3.2] to the case where the operators $B$ and $C$ occurring in the theorem have closed range. Note that when $T=I$ and $S=I$, we recover the Eckhart–Young theorem.
@@
[5, Theorem 3.2, Remark 3.5] Let $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3},\mathcal{H}_{4}$ be Hilbert spaces and let $T\in\mathcal{B}(\mathcal{H}_{3},\mathcal{H}_{4})$, $S\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$ both have closed range and let $M\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Suppose $P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}}$ has nonincreasing singular value sequence $(\sigma_{i})_{i}\in\ell^{2}([0,\infty))$. Then, for each rank-$r$ truncated SVD $(P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}})_{r}$ of $P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}}$,
@@
Theorem 5.10 and Theorem 5.11 generalise the results of [33, Theorem 4.1 and Theorem 4.2] to an infinite-dimensional setting, and add a uniqueness statement. We note that in both considered approximation classes $\mathscr{M}_{r}^{(i)}$, $i\in\{1,2\}$, the optimal operator $A^{\textup{opt},(i)}_{r}$ maps into $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$, just like the exact operator $\mathcal{C}_{\textup{pos}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}$ in (3a).

### v3

@@ 3 Equivalent Gaussian measures and Bayesian inference
We quantify posterior approximation errors using various divergences. Let $\nu_{2}$ be a target measure on $\mathcal{H}$ and $\nu_{1}$ an approximation of $\nu_{2}$ satisfying $\nu_{1}\sim\nu_{2}$. We use the $\rho$-Rényi divergence, the forward Kullback-Leibler (KL) divergence, the Amari $\alpha$-divergence for $\alpha\in(0,1)$ and the Hellinger distance, defined respectively by,
@@
We refer to $D_{\textup{KL}}(\nu_{1}\|\nu_{2})$ as the ‘reverse KL divergence’. We do not distinguish between forward Rényi divergences $D_{\textup{Ren},\rho}(\nu_{2}\|\nu_{1})$ and reverse Rényi divergences $D_{\textup{Ren},\rho}(\nu_{1}\|\nu_{2})$, because of the ‘skew symmetry’ of the Rényi divergence: $D_{\textup{Ren},\rho}(\nu_{1}\|\nu_{2})=D_{\textup{Ren},1-\rho}(\nu_{2}\|\nu_{1})$, c.f. [42, Proposition 2].
@@
where (6) follows by [28, eqs. (134)–(135)]. It is then straightforward to show, c.f. [10, Remarks 3.10 and 3.11] that minimising the Amari-$\alpha$ divergence over $\nu_{1}$ is equivalent to minimising the $\alpha$-Rényi divergence over $\nu_{1}$. Furthermore, minimising the Hellinger distance over $\nu_{1}$ is equivalent to minimising the $\frac{1}{2}$-Rényi divergence over $\nu_{1}$. The divergence $D_{\textup{Ren},\frac{1}{2}}$ is also known as the Bhattacharyya distance, and is a metric.
@@
For a proof, see e.g. [4, Corollary 6.4.11] or [17, Theorem 2.25]. For injective covariances $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ such that items (i) and (iii) in Theorem 3.2 hold, we define
@@
If $\nu_{1}$ and $\nu_{2}$ are Gaussian measures, then the above divergences can be expressed explicitly in terms of the means and covariances of $\nu_{1}$ and $\nu_{2}$ using $R(\cdot\|\cdot)$ defined in (7). These formulations rely on a generalisation of the determinant to infinite-dimensional Hilbert spaces. For $A\in L_{1}(\mathcal{H})$, the so-called ‘Fredholm determinant’ $\det(I+A)$ can be defined, and if only $A\in L_{2}(\mathcal{H})$, then the notion of ‘Hilbert–Carleman determinant’ $\det_{2}(I+A)$ can be used. The Fredholm and Hilbert–Carleman determinants are defined on respectively trace-class and Hilbert–Schmidt perturbations of the identity. In finite dimensions, every operator is a trace-class and Hilbert–Schmidt perturbation of the identity, and hence these generalised determinants are defined everywhere in this case. In fact, the Fredholm determinant agrees with the standard determinant in this case. We refer to [35, Theorem 3.2, Theorem 6.2] or [36, Lemma 3.3, Theorem 9.2] for details.
The following result holds when $\mathcal{H}$ is a separable Hilbert space of finite or infinite dimension. The proof is a direct application of [28, Theorems 14 and 15].
@@
[10, Theorem 3.8] Let $m_{1},m_{2}\in\mathcal{H}$ and $\mathcal{C}_{1},\mathcal{C}_{2}\in L_{2}(\mathcal{H})_{\mathbb{R}}$ be positive. If $\mathcal{N}(m_{1},\mathcal{C}_{1})\sim\mathcal{N}(m_{2},\mathcal{C}_{2})$, then
@@
The prior and posterior distributions in (1) are equivalent, for every realisation $y$ in a set of probability 1, c.f. [40, Theorem 6.23]. The Hessian $H$ defined in (2) has rank $n$, hence the posterior precision is a finite-rank update of the prior by (3c). Using the operators $R(\mathcal{C}_{\textup{pr}}\|\mathcal{C}_{\textup{pos}})$ and $R(\mathcal{C}_{\textup{pos}}\|\mathcal{C}_{\textup{pr}})$ and Theorem 3.2, we can obtain the following relations between the prior-preconditioned Hessian $\mathcal{C}_{\textup{pr}}^{1/2}H\mathcal{C}_{\textup{pr}}^{1/2}$ in (9a), the posterior-preconditioned Hessian in (9b), and the ‘pencil’ defined by the prior and the posterior covariance in (9c). The prior-preconditioned Hessian combines prior covariance information with information contained in the Hessian, i.e. information on the forward map, noise covariance, and data dimension. Recall the notation $v\otimes w$ for $u,w\in\mathcal{H}$ from Section 1.4.
@@
[10, Proposition 3.7] There exists a nondecreasing sequence $(\lambda_{i})_{i}\in\ell^{2}((-1,0])$ consisting of exactly $\operatorname{rank}\left({H}\right)$ nonzero elements and ONBs $(w_{i})_{i}$ and $(v_{i})_{i}$ of $\mathcal{H}$ such that $w_{i},v_{i}\in\operatorname{ran}{\mathcal{C}_{\textup{pr}}^{1/2}}$ and $v_{i}=\sqrt{1+\lambda_{i}}\mathcal{C}_{\textup{pos}}^{-1/2}\mathcal{C}_{\textup{pr}}^{1/2}w_{i}$ for every $i\in\mathbb{N}$, and
@@
We note that the eigenvalues $(\frac{-\lambda_{i}}{1+\lambda_{i}})_{i}$ of (9a) relate to the eigenvalues $(\delta_{i}^{2})_{i}$ of [39, eq. (2.8)] via the transformation $\lambda_{i}=\eta(\delta_{i}^{2})$, $\delta_{i}^{2}=\eta(\lambda_{i})$ with $\eta(x)=\frac{-x}{1+x}$ for $x\in(-1,\infty)$.
@@
The significance of (10) is that the posterior variance along the span of $\mathcal{C}_{\textup{pr}}^{-1/2}w_{i}$ is smaller than the prior variance along the same subspace by a factor of $(1+\tfrac{-\lambda_{i}}{1+\lambda_{i}})^{-1}$, for $i\in\mathbb{N}$. This was observed in the finite-dimensional case in [39, eq. (3.4)]. Thus, Proposition 3.4 implies that finite-dimensional data can only inform finitely many directions in parameter space, in the sense that posterior variance is reduced relative to prior variance only over a finite-dimensional subspace. The directions $(\mathcal{C}_{\textup{pr}}^{-1/2}w_{i})_{i\leq\operatorname{rank}\left({H}\right)}$ are orthogonal with respect to the $\mathcal{C}_{\textup{pr}}$-weighted inner product $\langle h_{1},h_{2}\rangle_{\mathcal{C}_{\textup{pr}}}\coloneqq\langle\mathcal{C}_{\textup{pr}}h_{1},h_{2}\rangle$, and not the unweighted inner product of $\mathcal{H}$.
@@
are subject to the largest relative variance reduction possible. Since $\mathcal{C}_{\textup{pr}}^{1/2}$ is injective, we thus conclude the following: among all $r$-dimensional subspaces of $\mathcal{H}$, it is the $r$-dimensional subspace $W_{r}$ that contains those $r$ linearly-independent directions in which the relative variance reduction is largest. This generalises the conclusion of [39, Section 3.1] to infinite dimensions.
@@ 4 Optimal approximation of the covariance
This section discusses low-rank posterior covariance approximation, using [10, Theorem 4.21]. This approximation serves as a basis for the joint mean and covariance approximation discussed in Section 6.
@@
Since $\mathcal{C}_{\textup{pr}}-KK^{*}\in\mathscr{C}_{r}$ is positive and self-adjoint, it is an injective covariance operator. Furthermore, it is stated in [10, Corollary 4.9] that for every $\mathcal{C}\in\mathscr{C}_{r}$ it holds that $\mathcal{N}(m_{\textup{pos}}(y),\mathcal{C})$ is equivalent to the exact posterior. Since $\mathcal{C}_{\textup{pos}}$ does not depend on $y$, this equivalence holds for all $y$ simultaneously. This equivalence holds because of the range condition $\operatorname{ran}{K}\subset\operatorname{ran}{\mathcal{C}_{\textup{pr}}}$. Furthermore, the assumption $K\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ implies the rank restriction $\operatorname{rank}\left({K}\right)\leq r$. Thus, for $r$ small compared to $n$, $\mathcal{C}_{\textup{pr}}-KK^{*}$ can be interpreted as a low-rank update of $\mathcal{C}_{\textup{pr}}.$ Therefore, the class $\mathscr{C}_{r}$ provides an extension to infinite dimensions of the finite-dimensional updates considered in [39].
@@
The solution to Problem 4.1 now follows directly from [10, Lemma 4.2(iii)] and from [10, Theorem 4.21] applied with $f(x)\leftarrow f_{\textup{KL}}(\tfrac{-x}{1+x})$ , where
@@
Note that the results of [10, Theorem 4.21 and Corollary 4.23] are more general. They state that $\mathcal{C}^{\textup{opt}}_{r}$ is not only the optimal low-rank approximation of $\mathcal{C}_{\textup{pos}}$ for the reverse KL divergence, but simultaneously also for all divergences in a more general class of divergences, including the forward KL divergence, the Hellinger distance, the Rényi divergences and the Amari $\alpha$-divergences for $\alpha\in(0,1)$.
@@ 5 Optimal approximation of the mean
We shall assess the approximation quality of an approximate posterior mean by averaging the mean-dependent term for the Rényi divergence and the forward and reverse KL divergence over all possible realisations $y$ of $Y$. By averaging over $Y$, the optimal operator $A$ will be data-independent, i.e. will not depend on a specific realisation $y$ of $Y$. While averaging over $Y$ implies that the resulting posterior mean approximations are not optimal in general for a specific realisation $y$ of $Y$, this approach has the benefit that $A$ can be constructed before observing the data. This leads to an offline-online approach to posterior mean approximation: the preliminary ‘offline’ stage computes one operator, which can then can be applied in the subsequent ‘online’ stage to any realisation of the data. This is in analogy to the finite-dimensional case studied in [39, Section 4.1] and its generalisation to certain nonlinear forward models and to losses with respect to the average Amari $\alpha$-divergences as studied in [27, Section 5]. Furthermore, averaging over $Y$ enables us to exploit recent work on reduced-rank operator approximation [9].
@@
The Bayes risk $\mathcal{R}(A)\coloneqq\mathbb{E}\left[\left\lVert AY-X\right\rVert_{\mathcal{C}_{\textup{pos}}^{-1}}^{2}\right]$ for $A\in\mathscr{M}_{r}^{(i)}$, $i=1,2$, considered in [39, Section 4.1] is not well-defined, since the event $\{X\in\operatorname{dom}{\mathcal{C}_{\textup{pos}}^{-1/2}}\}$ occurs with probability 0. However, one can show that $\mathcal{R}(A)=\mathbb{E}\left[\left\lVert AY-m_{\textup{pos}}(Y)\right\rVert_{\mathcal{C}_{\textup{pos}}^{-1}}^{2}\right]+\dim{\mathcal{H}}$ if $\dim{\mathcal{H}}<\infty$. Thus, not only does the approximation error (18) used in Problem 5.1 have a natural interpretation as the mean-dependent term of the Rényi, Amari, forward and reverse KL divergences, it also captures the relevant contribution to the Bayes risk which involves the approximation.
@@
$\begin{split}\widetilde{\mathscr{M}}^{(1)}_{r}&\coloneqq\{(S_{\textup{pos}}^{-1}\mathcal{C}_{\textup{pr}}-\widetilde{B})G^{*}\mathcal{C}_{\textup{obs}}^{-1}:\ \widetilde{B}\in\mathcal{B}_{00,r}(\mathcal{H})\},\\ \widetilde{\mathscr{M}}^{(2)}_{r}&\coloneqq\mathcal{B}_{00,r}(\mathbb{R}^{n},\mathcal{H}).\end{split}$ (23)
@@
It is shown in items (iii) and (iv) of the following result that Problem 5.4 is a reformulation of Problem 5.1. Using Theorem 3.2, item (i) of the following result also provides an explicit description of the approximation classes $\mathscr{M}^{(i)}_{r}$ of (4) in terms of the ranges of the operators $A$ and $B$, while item (ii) relates these classes to the classes $\widetilde{\mathscr{M}}^{(i)}_{r}$ from (23).
@@
Let $r\leq n$ and $i=1,2$. Let $S_{\textup{pos}}$ be as defined in (21), let $\mathscr{M}^{(i)}_{r}$ be as in (4) and let $\widetilde{\mathscr{M}}^{{(i)}}_{r}$ be as in (23). Then,
@@
    $\widetilde{\mathscr{M}}^{(i)}_{r}=S_{\textup{pos}}^{-1}\mathscr{M}^{(i)}_{r}$,
@@
In order to solve Problem 5.4, we use a result on reduced-rank operator approximation in $L_{2}(\mathcal{H})$ norm, proven in [9]. It is a generalised version of the Eckhart–Young theorem. Recall that compact operators, in particular Hilbert–Schmidt operators and finite-rank operators, have an SVD, c.f. Lemma A.3. Also recall the definition of the Moore–Penrose inverse $C^{\dagger}$ of $C\in\mathcal{B}(\mathcal{H})$ from Section 1.4. If $C$ has closed range, then $C^{\dagger}$ is bounded, c.f. [18, Proposition 2.4]. The following is an application of [9, Theorem 3.2] to the case where the operators $B$ and $C$ occurring in the theorem have closed range. Note that when $T=I$ and $S=I$, we recover the Eckhart–Young theorem.
@@
[9, Theorem 3.2, Remark 3.5] Let $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}_{3},\mathcal{H}_{4}$ be Hilbert spaces and let $T\in\mathcal{B}(\mathcal{H}_{3},\mathcal{H}_{4})$, $S\in\mathcal{B}(\mathcal{H}_{1},\mathcal{H}_{2})$ both have closed range and let $M\in L_{2}(\mathcal{H}_{1},\mathcal{H}_{4})$. Suppose $P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}}$ has nonincreasing singular value sequence $(\sigma_{i})_{i}\in\ell^{2}([0,\infty))$. Then, for each rank-$r$ truncated SVD $(P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}})_{r}$ of $P_{\operatorname{ran}{T}}MP_{\ker{S}^{\perp}}$,
@@
Theorem 5.10 and Theorem 5.11 generalise the results of [39, Theorem 4.1 and Theorem 4.2] to an infinite-dimensional setting, and add a uniqueness statement. We note that in both considered approximation classes $\mathscr{M}_{r}^{(i)}$, $i\in\{1,2\}$, the optimal operator $A^{\textup{opt},(i)}_{r}$ maps into $\operatorname{ran}{\mathcal{C}_{\textup{pos}}}$, just like the exact operator $\mathcal{C}_{\textup{pos}}G^{*}\mathcal{C}_{\textup{obs}}^{-1}$ in (3a).